OpenStudy (anonymous):

A student prepared a sample of aspirin with a mass of 3.92g. If the reaction produce 80.00% yield, what was the theorectical yield of the reactant?

OpenStudy (xishem):

We can set up a proportion here... \[\frac{yield_{actual}}{yield_{theoretical}}=\frac{mass_{actual}}{mass_{theoretical}}\]Rewritten in terms of the theoretical mass...\[mass_{theoretical}=\frac{(mass_{actual})(yield_{theoretical})}{yield_{actual}}\]Since, we expect to get 100% from such processes, theoretically...\[mass_{theoretical}=\frac{(3.92g)(100\%)}{80\%}=4.90g\]

OpenStudy (anonymous):

thanks